OpenStudy (anonymous):

Write an equation in slope intercept form perpendicular to y=-3x+1 passes through 2,2

OpenStudy (anonymous):

I came up with y=3x-4

HanAkoSolo (jamierox4ev3r):

right, but when the equation is perpendicular, you use the reciprocal. THat would mean that -3 would become \(\Large\frac{1}{3}\)

OpenStudy (anonymous):

OK so that means y=1/3x-4

OpenStudy (anonymous):

?

HanAkoSolo (jamierox4ev3r):

lol sorry i was afk, but that would actually change your y-intercept. you need to calculate based on the new y-intercept. You equation would look like this 2=\(\frac{1}{3}\)(2)+b

HanAkoSolo (jamierox4ev3r):

then solve for b and that would be your y-intercept. Can you do that for me @JoStudy?